# gPC-based robustness analysis of neural systems through probabilistic recurrence metrics

### v1

@@ 1 Introduction
In systems biology, robustness refers to the ability of a system to preserve a qualitative property despite uncertainties, variability or environmental fluctuations [kitano2004biological, Khammash2016, BG2021structural, PKG2025]. In particular, systems in neuroscience can robustly preserve complex signalling patterns, often called regimes, both at the single neuron level and at the neural population level [bray2024adaptive, mirzakhalili2017synaptic, li2016robust, aerts2016brain, maclean2003activity], despite being subject to environmental disturbances and intrinsic uncertainties [koch2004biophysics, faisal2008noise]. Understanding the mechanisms that guarantee such robustness, as well as quantifying the uncertainty levels that are compatible with preservation of a given life-sustaining signalling pattern, enables a better interpretation of recorded signals and a deeper insight into neural dynamics and neural diseases. The study of dynamical regimes and of their robustness in neural models also supports the robust design of artificial systems inspired by dynamics in neuroscience [hassabis2017neuroscience], as well as future biomedical applications and medical interventions, e.g., to monitor depth of anaesthesia [kuhlmann2016neural], treat Parkinson’s disease [kerr2013cortical] and epilepsy [subramaniyam2024sensitivity], and contrast epileptic seizures [ashourvan2020model].
@@
The analysis of interpretable dynamical models of neural activity is crucial to understand neural behaviours [izhikevich2007dynamical]. To assess how varying key parameters, or combinations thereof, affects the behaviour of neural systems and the resulting (number and type of) observed regimes (such as equilibria), existing works usually employ deterministic methods, such as sensitivity analysis, parameter sweeping or bifurcation analysis [barrio2011parameter, ashwin2016mathematical], where parameter values are assumed to be constant and deterministic. These methods to assess the number and properties of dynamical regimes in a parameter space, and the regime changes associated with parameter variations, do not account for the possibility that biological parameters are affected by probabilistic uncertainties. Also, given experimental data, traditional modelling typically employs deterministic methods to identify the single best-fit value of the system parameters that matches the observed biological phenomenon and then uses the resulting model for prediction; the usefulness and reliability of the obtained predictions critically depend on the accuracy, precision and completeness of measurements [gutenkunst2007universally]. However, parametric uncertainty and variability strongly affect real-life neuronal systems, and thus the ensuing measurements, as observed in experimental settings [goaillard2021ion]. Probabilistic parametric uncertainty causes the system parameters to take values according to a probability distribution, within feasible bounds including a nominal value, thus yielding different system realisations (such as those observed in distinct neurons within a population) that may exhibit the same qualitative physical behaviour despite having different system parameters. This exactly matches real-life scenarios: different cells, neurons and structures with the same function do not typically share the same parameter values, since these are associated with physiological processes, such as ion fluxes through membrane channels, that are subject to variability and thus induce heterogeneity, across cells or over time [nusser2009variability, goaillard2021ion]. Uncertainty and variability, therefore, are intrinsic and cannot be neglected in the study of neuronal systems.
@@
Mathematically, uncertainty can be characterised either in a deterministic or in a probabilistic setting. Studying the effect of deterministic parametric uncertainties is the goal of classic robustness analysis, which has a long tradition in control theory [barmish1994new]. Robustness analysis is agnostic about the exact parameter values realised in practice, and provides a “yes” or “no” answer to the question whether a property of interest (in our case, a regime) is preserved robustly, i.e., for all possible parameter realisations within a known bounding set in the parameter space. The resulting sufficient conditions for the desired system performance are often overly conservative [barmish1994new, Chapter 14] and have an “all-or-nothing” nature: if, even for a single parameter choice within the bounding set, the property does not hold, then robustness cannot be guaranteed. In nature, however, certain proportions of defective cells and processes may be present while proper biological functions are maintained nonetheless [mojtahedi2016cell]. Hence, preservation of a life-sustaining behaviour “on average” (e.g., through a critical mass of properly functioning cells) is what truly matters. This observation fosters the need to investigate probabilistic parameter variations in biological systems, and their effects on the preservation of biological functions and regimes.
@@
In this work, to study the robustness of neurological systems with respect to parametric uncertainty, we resort to probabilistic robustness analysis (PRA) [tempo2013randomized]. Differently from deterministic methods such as sensitivity and bifurcation analysis, and from classic parametric robustness analysis, PRA allows us to consider probabilistic parametric uncertainty in neurological models and more accurately analyse the effects of randomness in parameter realisations on the resulting (stochastic) behaviour of neural systems. In fact, PRA of a system whose uncertain parameters can vary, within a known bounding set in the parameter space, according to a known probability distribution (which encodes likelihood belief, estimation error, or population variability) has several advantageous properties. (i) PRA identifies uncertainty regions in the parameter space where the same qualitative system behaviour is preserved [goldman2001global, foster1993significance, beer1999evolution, marder2011variability] in probability. PRA can not only assess whether a system preserves a regime of interest in expectation in the presence of given parametric uncertainties affecting its nominal parameters, but also quantify how much parametric uncertainty affecting its nominal parameters a system can withstand, while still preserving a regime of interest in expectation. (ii) PRA helps interpret the emergent average effects of probabilistic parametric uncertainty: it allows to quantify, predict and analyse the mean neural activity of ensembles of heterogeneous neurons, which is considered to be a more robust and reliable proxy for neural functions than the activity of single neurons, individually more fragile and prone to malfunctioning [pouget2000information]. Importantly, the signal associated with nominal deterministic values of the parameters may be completely different from the mean signal realisation obtained when the parameters are subject to parametric uncertainty within a parameter bounding set that contains the nominal parameter values (see Supplementary Figure S4); hence, for instance, regime identification is much more difficult, compared to the deterministic setting. If all parameter realisations in the bounding set yield the same qualitative behaviour (as in deterministic robustness analysis), then the mean realisation in the set is more likely, but is not guaranteed, to preserve the same qualitative behaviour. These observations make the question of mean regime preservation under probabilistic parametric uncertainty much more intricate than its deterministic robustness counterpart. In particular, studying the deterministic realisation corresponding to the average of multiple values of the system parameters cannot provide any meaningful information on the mean system realisation. Indeed, for systems in neuroscience it is well known that the signal associated with the mean of multiple parameter values may not be representative of the regime associated with each of the original parameter values [golowasch2002failure] (see, e.g., Supplementary Figure S3). Systematic PRA tackles this complexity. (iii) When comparing models to data, in-vivo recording methods such as electro-encephalograms (EEG) or local field potentials (LFP) measure the mean signals over batches of neurons [perinelli2020chasing, lucas2024topological]. Therefore, PRA is a promising tool to understand whether regimes obtained from single-neuron models can robustly predict observed ensemble recordings, and to improve the synergy between mathematical models and real-life data through more reliable parameter fitting. (iv) The ability to detect models that reproduce observed neurological phenomena only for specific parameter choices and cannot accommodate even reasonably small parameter variations while preserving the intended neurological behaviour, as opposed to models that display robust neurologically-meaningful regimes in the face of natural parametric uncertainty, makes PRA precious to falsify models and guide model selection, also in combination with suitably designed experiments [goldman2001global].
@@
In this paper, we develop a novel methodology for the efficient and systematic PRA of uncertain neuronal systems and we demonstrate its efficacy on well-established dynamical models from mathematical neuroscience: the Hindmarsh-Rose (HR) model [hindmarsh1984model] and the Jansen-Rit (JR) model [jansen1995electroencephalogram]. We consider a class of models of the form $\dot{\mathbf{x}}=f(\mathbf{x},Z)$ subject to probabilistic parametric uncertainty, where the parameters in $Z$ are independent random variables characterised by a known uniform probability distribution, supported on a bounded set that contains the nominal parameter values. Then, the state $\mathbf{x}(t;Z)$ is a stochastic processes and each realisation of the system trajectory depends on a realisation of the parameters in $Z$. In this probabilistic framework, we aim to identify regions in the parameter space where the regime associated with the deterministic nominal model is preserved in expectation despite parametric uncertainties, so as to distinguish between regimes and quantify the largest uncertainty level compatible with probabilistic robustness of a regime of interest. To assess probabilistic robustness, we focus on time-averages (means) of output signals of interest (obtained from the imposed parameter distributions); to efficiently compute mean neural activities under parametric uncertainty, we resort to the generalised polynomial chaos (gPC) method [xiu2003modeling], which has two major advantages: first, it effectively incorporates parametric uncertainty into the analysed models and provides explicit analytic expressions that directly capture the uncertainty effects; second, it provides a favourable trade-off between accuracy and computational costs [sutulovic2024efficient], often being much more efficient than popular Monte Carlo methods, a commonly used alternative for model-based PRA. Then, we propose a computational methodology to extract patterns from the obtained mean time-series through the generation of recurrence plots [eckmann1995recurrence]. The identification of distinct geometric patterns within recurrence plots enables to distinguish between regimes; we show that our proposed approach, based on blob counts, successfully copes with the presence of parametric uncertainty (while other metrics often used for regime identification in deterministic models, such as inter-spike intervals, duty cycles or spike counts, fail to identify regimes and to infer reliable information about their preservation). Finally, our methodology produces probabilistic regime preservation (PRP) plots that contain essential information on the dynamical regimes induced by the system under study, and their robustness to parametric uncertainties, in a parameter space of interest. Differently from classical deterministic methods, such as bifurcation plots, the obtained PRP plots offer a probabilistic perspective on the mean behaviour of the system under investigation, and on the preservation of its regimes when the system parameters are affected by probabilistic uncertainty. We present in detail our proposed computational methodology for PRA of neural systems in Section 2, after providing some necessary background on the considered models and on the gPC technique. Then, in Section 3, we apply our new approach to the HR and the JR models, and discuss the significance of the obtained PRP results for their robustness to parametric uncertainty. A concluding discussion is given in Section 4. Additional information and simulations are provided in the Supplementary Material.
@@ 2 Preservation of regimes under parametric uncertainty
Figure 1: Workflow of our methodology for probabilistic robustness analysis (PRA) described in Section 2. (a) Consider a dynamical system of interest, where the evolution of the state vector $\mathbf{x}\in\mathbb{R}^{n}$ depends on a vector $Z$ of $d$ uncertain parameters, independent and distributed according to uniform distributions within known bounding intervals. (b) Derive a surrogate model of the system with the gPC method. (c) Based on the surrogate model, efficiently compute the mean system output (blue) with respect to the parametric uncertainty $Z$; the corresponding deterministic output of the nominal, uncertainty-free system is shown (purple) for comparison. (d) To unravel key geometric patterns within the time series associated with the mean system output, construct the corresponding recurrence plot (left). Then, automatically extract persistent features (blob count) of recurrence plots corresponding to increasing uncertainty levels, so as to assess regime preservation: the regime is preserved for all the uncertainty levels that achieve a similar blob count, while it is no longer preserved once the blob count drops (right). (e) Capture the overall information about dynamical regimes and their robustness to parametric uncertainty in a probabilistic regime preservation (PRP) plot, where the colour encodes blob counts and the length of each side of the rectangles quantifies how much uncertainty the system can withstand in the corresponding parameter value while maintaining the same regime in expectation. (f) For comparison, in the same parameter space as the PRP plot, we present a bifurcation plot (image courtesy of Springer, see [barrio2011parameter] for additional details) that identifies the regimes resulting from different deterministic values of the parameters.
@@ 2.1 Class of considered systems subject to parametric uncertainties
We consider models from mathematical neuroscience that are described by an ordinary differential equation (ODE) system of the form $\dot{\mathbf{x}}=f(\mathbf{x},Z)$, where $\mathbf{x}\in\mathbb{R}^{n}$ is the state vector and the uncertain system parameters of interest are stacked in a parametric uncertainty vector $Z\in\mathbb{R}^{d}$. We assume that the components of $Z$ are independent random variables, characterised by a known probability distribution supported on a bounded set representing the parameter space of interest for the considered model. Many different distributions of $Z$ may be handled by our proposed approach [xiu2003modeling]. Since the uniform distribution has been observed to be the worst-case distribution for PRA [BARMISH!!], and we can theoretically transform any continuous random variable into a uniformly distributed one through a quantile transformation [kallenberg2002foundations], in the absence of additional information on $Z$, we assume that each component $Z_{i}$, $i=1,\ldots,d$, is uniformly distributed over an uncertainty interval containing the nominal value $Z_{i}^{*}$. Precisely, we therefore have $Z_{i}\sim\mathcal{U}([Z_{i,{\min}},Z_{i,{\max}}])$, with $Z_{i,{\min}}\leq Z_{i}^{*}\leq Z_{i,{\max}}$, $i=1,\ldots,d$. The resulting ODE model subject to parametric uncertainty is
@@
where $f\colon\mathbb{R}^{n}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$, the hyper-rectangle $\mathcal{A}$ is the uncertainty bounding set and $\mathbf{x}(t;Z)$ is a stochastic process (Figure 1a); in this probabilistic setting, different system evolutions are possible depending on the specific realisation of the parameters in $Z$. In (1), stochasticity arises from probabilistic uncertainty in the system parameters, and not from a stochastic process driving the dynamics as it happens for instance in stochastic differential equations [oksendal2013stochastic].
@@
To demonstrate our proposed methodology, we focus on two celebrated models from neuroscience: the Hindmarsh-Rose (HR) model for single-neuron dynamics [hindmarsh1984model] and the neural mass mean-field Jansen-Rit model for a single cortical column [jansen1995electroencephalogram].
The HR model [hindmarsh1984model] is widely used to reproduce the spiking and bursting dynamics of single neurons observed in experimental data [gu2013biological]; however, its robustness to probabilistic parametric uncertainty has not been studied so far. In the adimensionalised ODEs of the HR model
@@
variable $x_{1}(t)$ represents the neuron membrane potential, $x_{2}(t)$ a fast recovery current, and $x_{3}(t)$ a slow adaptation current that enables bursting (i.e., trains of multiple membrane potential spikes). $I$ is an injected current, $x_{R}$ is the reference resting membrane potential, $r$ is the timescale ratio between fast and slow currents, while parameters $a$, $b$, $c$, $d$, $s$ are related to the system nullclines [hindmarsh1984model]. We consider $Z_{\text{HR}}=[b,I]^{\top}$ as the parameters of interest, and set the other parameters to the standard values in [hindmarsh1984model]: $a=1$, $c=1$, $d=5$, $s=4$, $x_{R}=-8/5$, $r=0.01$. In a deterministic setting, for different values of the system parameters, the state trajectories $\mathbf{x}(t)=[x_{1}(t),x_{2}(t),x_{3}(t)]^{\top}$, $t\geq 0$, of this relatively simple model can exhibit multiple qualitatively distinct behaviours, which we call dynamical regimes. We focus on the membrane potential, $y(t)=x_{1}(t)$, as an output signal of interest; examples of realisations of $x_{1}(t)$ for different parameter values that pertain to different dynamical regimes, ranging from quiescence to bursting, are shown in Figure 2 (top).
The JR model [jansen1995electroencephalogram] is a neural mass mean-field model that describes, rather than the dynamics of individual neurons, the generation of EEG signals by a single cortical column. In particular, the cortical column is represented through the interaction of the mean-field activity of three neuronal populations: pyramidal neurons, excitatory interneurons and inhibitory interneurons. In the corresponding ODE system
@@
the states $x_{i}(t)$, $i=1,2,3$, are neural potentials ($x_{1}$ is the mean excitatory postsynaptic potential, $x_{2}$ is the mean membrane potential of pyramidal cells, $x_{3}$ is the mean inhibitory postsynaptic potential), while $S(V)=\nu_{M}/\left(1+e^{r(V_{0}-V)}\right)$ is a sigmoidal function, where $\nu_{M}$ is the maximal firing rate of the family of neurons, $V_{0}$ is the excitability threshold of the population, $r$ is the firing threshold variance (i.e., the slope of the sigmoid at $V_{0}$). Parameters $A$ and $B$ are the maximum amplitude of excitatory and inhibitory postsynaptic potentials, respectively, while $a$ and $b$ are lumped parameters related to time constants in the dendritic network for excitatory and inhibitory synapses, respectively. All connectivity constants $C_{i}$, describing the strength of the links between pyramidal neurons and inhibitory and excitatory interneurons, are proportional to a single parameter $C$: $C_{1}=C$, $C_{2}=0.8C$, $C_{3}=0.25C$, $C_{4}=0.25C$, with $C=135$ [jansen1995electroencephalogram]. We investigate $Z_{\text{JR}}=[A,B]$ as the parameters of interest, for varying values of $p$ and $C$, while all other parameters are fixed as in [jansen1995electroencephalogram]: $a=100$, $b=50$, $V_{0}=6$, $\nu_{M}=5$, $r=0.56$. In a deterministic setting, the EEG-like output signal $y(t)=x_{2}(t)-x_{3}(t)$ exhibits different qualitative dynamical behaviours (regimes) as shown in Supplementary Figure S2.
@@
We consider the HR parameter space $b_{\min},b_{\max}]\times[I_{\min},I_{\max}]=[2.5,3.3]\times[2.2,4.4]$. Figure 1f shows the corresponding bifurcation diagram, from [barrio2011parameter], in the deterministic setting: regions with different colours correspond to deterministic parameter values that result in regimes with different characteristics (exemplified in Figure 2). In our numerical experiments, we choose nominal parameter values $b^{*},I^{*}]^{\top}$ that lie within the interior of a region associated with one of the deterministic nominal regimes in Figure 1f. Probabilistic parametric uncertainty affecting the nominal parameter values may lead to loss of regime identifiability from the mean output trajectory realisation (see Figure 1c and Supplementary Figure S4 for examples). Our goal is thus to develop a computational framework that allows us to detect regimes in the presence of parametric uncertainty and identify the highest possible uncertainty level (corresponding to the largest possible neighbourhood of the nominal parameters, given predefined specifications) where the nominal regime, exhibited by the system output under the deterministic nominal parameters, is preserved in expectation. To this aim, we design dedicated metrics and apply them to the mean output signal, so as to relate it to the nominal regime.

### v2

@@ 1 Introduction
In systems biology, robustness refers to the ability of a system to preserve a qualitative property despite uncertainties, variability or environmental fluctuations [35, 34, 10, 59]. In particular, systems in neuroscience can robustly preserve complex signalling patterns, often called regimes, both at the single neuron level and at the neural population level [12, 48, 40, 1, 44], despite being subject to environmental disturbances and intrinsic uncertainties [36, 16]. Understanding the mechanisms that guarantee such robustness, as well as quantifying the uncertainty levels that are compatible with preservation of a given life-sustaining signalling pattern, enables a better interpretation of recorded signals and a deeper insight into neural dynamics and neural diseases. The study of dynamical regimes and of their robustness in neural models also supports the robust design of artificial systems inspired by dynamics in neuroscience [25], as well as future biomedical applications and medical interventions, e.g., to monitor depth of anaesthesia [38], treat Parkinson’s disease [33] and epilepsy [67], and contrast epileptic seizures [3].
@@
The analysis of interpretable dynamical models of neural activity is crucial to understand neural behaviours [29]. To assess how varying key parameters, or combinations thereof, affects the behaviour of neural systems and the resulting (number and type of) observed regimes (such as equilibria), existing works usually employ deterministic methods, such as sensitivity analysis, parameter sweeping or bifurcation analysis [7, 4], where parameter values are assumed to be constant and deterministic. These methods to assess the number and properties of dynamical regimes in a parameter space, and the regime changes associated with parameter variations, do not account for the possibility that biological parameters are affected by probabilistic uncertainties. Also, given experimental data, traditional modelling typically employs deterministic methods to identify the single best-fit value of the system parameters that matches the observed biological phenomenon and then uses the resulting model for prediction; the usefulness and reliability of the obtained predictions critically depend on the accuracy, precision and completeness of measurements [24]. However, parametric uncertainty and variability strongly affect real-life neuronal systems, and thus the ensuing measurements, as observed in experimental settings [20]. Probabilistic parametric uncertainty causes the system parameters to take values according to a probability distribution, within feasible bounds including a nominal value, thus yielding different system realisations (such as those observed in distinct neurons within a population) that may exhibit the same qualitative physical behaviour despite having different system parameters. This exactly matches real-life scenarios: different cells, neurons and structures with the same function do not typically share the same parameter values, since these are associated with physiological processes, such as ion fluxes through membrane channels, that are subject to variability and thus induce heterogeneity, across cells or over time [50, 20]. Uncertainty and variability, therefore, are intrinsic and cannot be neglected in the study of neuronal systems.
@@
Mathematically, uncertainty can be characterised either in a deterministic or in a probabilistic setting. Studying the effect of deterministic parametric uncertainties is the goal of classic robustness analysis, which has a long tradition in control theory [6]. Robustness analysis is agnostic about the exact parameter values realised in practice, and provides a “yes” or “no” answer to the question whether a property of interest (in our case, a regime) is preserved robustly, i.e., for all possible parameter realisations within a known bounding set in the parameter space. The resulting sufficient conditions for the desired system performance are often overly conservative [6, Chapter 14] and have an “all-or-nothing” nature: if, even for a single parameter choice within the bounding set, the property does not hold, then robustness cannot be guaranteed. In nature, however, certain proportions of defective cells and processes may be present while proper biological functions are maintained nonetheless [49]. Hence, preservation of a life-sustaining behaviour “on average” (e.g., through a critical mass of properly functioning cells) is what truly matters. This observation fosters the need to investigate probabilistic parameter variations in biological systems, and their effects on the preservation of biological functions and regimes.
@@
In this work, to study the robustness of neurological systems with respect to parametric uncertainty, we resort to probabilistic robustness analysis (PRA) [70]. Differently from deterministic methods such as sensitivity and bifurcation analysis, and from classic parametric robustness analysis, PRA allows us to consider probabilistic parametric uncertainty in neurological models and more accurately analyse the effects of randomness in parameter realisations on the resulting (stochastic) behaviour of neural systems. In fact, PRA of a system whose uncertain parameters can vary, within a known bounding set in the parameter space, according to a known probability distribution (which encodes likelihood belief, estimation error, or population variability) has several advantageous properties. (i) PRA identifies uncertainty regions in the parameter space where the same qualitative system behaviour is preserved [21, 17, 9, 45] in probability. PRA can not only assess whether a system preserves a regime of interest in expectation in the presence of given parametric uncertainties affecting its nominal parameters, but also quantify how much parametric uncertainty affecting its nominal parameters a system can withstand, while still preserving a regime of interest in expectation. (ii) PRA helps interpret the emergent average effects of probabilistic parametric uncertainty: it allows to quantify, predict and analyse the mean neural activity of ensembles of heterogeneous neurons, which is considered to be a more robust and reliable proxy for neural functions than the activity of single neurons, individually more fragile and prone to malfunctioning [57]. Importantly, the signal associated with nominal deterministic values of the parameters may be completely different from the mean signal realisation obtained when the parameters are subject to parametric uncertainty within a parameter bounding set that contains the nominal parameter values (see Supplementary Figure S4); hence, for instance, regime identification is much more difficult, compared to the deterministic setting. If all parameter realisations in the bounding set yield the same qualitative behaviour (as in deterministic robustness analysis), then the mean realisation in the set is more likely, but is not guaranteed, to preserve the same qualitative behaviour. These observations make the question of mean regime preservation under probabilistic parametric uncertainty much more intricate than its deterministic robustness counterpart. In particular, studying the deterministic realisation corresponding to the average of multiple values of the system parameters cannot provide any meaningful information on the mean system realisation. Indeed, for systems in neuroscience it is well known that the signal associated with the mean of multiple parameter values may not be representative of the regime associated with each of the original parameter values [22] (see, e.g., Supplementary Figure S3). Systematic PRA tackles this complexity. (iii) When comparing models to data, in-vivo recording methods such as electro-encephalograms (EEG) or local field potentials (LFP) measure the mean signals over batches of neurons [55, 43]. Therefore, PRA is a promising tool to understand whether regimes obtained from single-neuron models can robustly predict observed ensemble recordings, and to improve the synergy between mathematical models and real-life data through more reliable parameter fitting. (iv) The ability to detect models that reproduce observed neurological phenomena only for specific parameter choices and cannot accommodate even reasonably small parameter variations while preserving the intended neurological behaviour, as opposed to models that display robust neurologically-meaningful regimes in the face of natural parametric uncertainty, makes PRA precious to falsify models and guide model selection, also in combination with suitably designed experiments [21].
@@
In this paper, we develop a novel methodology for the efficient and systematic PRA of uncertain neuronal systems and we demonstrate its efficacy on well-established dynamical models from mathematical neuroscience: the Hindmarsh-Rose (HR) model [26] and the Jansen-Rit (JR) model [30]. We consider a class of models of the form $\dot{\mathbf{x}}=f(\mathbf{x},Z)$ subject to probabilistic parametric uncertainty, where the parameters in $Z$ are independent random variables characterised by a known uniform probability distribution, supported on a bounded set that contains the nominal parameter values. Then, the state $\mathbf{x}(t;Z)$ is a stochastic processes and each realisation of the system trajectory depends on a realisation of the parameters in $Z$. In this probabilistic framework, we aim to identify regions in the parameter space where the regime associated with the deterministic nominal model is preserved in expectation despite parametric uncertainties, so as to distinguish between regimes and quantify the largest uncertainty level compatible with probabilistic robustness of a regime of interest. To assess probabilistic robustness, we focus on time-averages (means) of output signals of interest (obtained from the imposed parameter distributions); to efficiently compute mean neural activities under parametric uncertainty, we resort to the generalised polynomial chaos (gPC) method [74], which has two major advantages: first, it effectively incorporates parametric uncertainty into the analysed models and provides explicit analytic expressions that directly capture the uncertainty effects; second, it provides a favourable trade-off between accuracy and computational costs [69], often being much more efficient than popular Monte Carlo methods, a commonly used alternative for model-based PRA. Then, we propose a computational methodology to extract patterns from the obtained mean time-series through the generation of recurrence plots [15]. The identification of distinct geometric patterns within recurrence plots enables to distinguish between regimes; we show that our proposed approach, based on blob counts, successfully copes with the presence of parametric uncertainty (while other metrics often used for regime identification in deterministic models, such as inter-spike intervals, duty cycles or spike counts, fail to identify regimes and to infer reliable information about their preservation). Finally, our methodology produces probabilistic regime preservation (PRP) plots that contain essential information on the dynamical regimes induced by the system under study, and their robustness to parametric uncertainties, in a parameter space of interest. Differently from classical deterministic methods, such as bifurcation plots, the obtained PRP plots offer a probabilistic perspective on the mean behaviour of the system under investigation, and on the preservation of its regimes when the system parameters are affected by probabilistic uncertainty. We present in detail our proposed computational methodology for PRA of neural systems in Section 2, after providing some necessary background on the considered models and on the gPC technique. Then, in Section 3, we apply our new approach to the HR and the JR models, and discuss the significance of the obtained PRP results for their robustness to parametric uncertainty. A concluding discussion is given in Section 4. Additional information and simulations are provided in the Supplementary Material.
@@ 2 Preservation of regimes under parametric uncertainty
Figure 1: Workflow of our methodology for probabilistic robustness analysis (PRA) described in Section 2. (a) Consider a dynamical system of interest, where the evolution of the state vector $\mathbf{x}\in\mathbb{R}^{n}$ depends on a vector $Z$ of $d$ uncertain parameters, independent and distributed according to uniform distributions within known bounding intervals. (b) Derive a surrogate model of the system with the gPC method. (c) Based on the surrogate model, efficiently compute the mean system output (blue) with respect to the parametric uncertainty $Z$; the corresponding deterministic output of the nominal, uncertainty-free system is shown (purple) for comparison. (d) To unravel key geometric patterns within the time series associated with the mean system output, construct the corresponding recurrence plot (left). Then, automatically extract persistent features (blob count) of recurrence plots corresponding to increasing uncertainty levels, so as to assess regime preservation: the regime is preserved for all the uncertainty levels that achieve a similar blob count, while it is no longer preserved once the blob count drops (right). (e) Capture the overall information about dynamical regimes and their robustness to parametric uncertainty in a probabilistic regime preservation (PRP) plot, where the colour encodes blob counts and the length of each side of the rectangles quantifies how much uncertainty the system can withstand in the corresponding parameter value while maintaining the same regime in expectation. (f) For comparison, in the same parameter space as the PRP plot, we present a bifurcation plot (image courtesy of Springer, see [7] for additional details) that identifies the regimes resulting from different deterministic values of the parameters.
@@ 2.1 Class of considered systems subject to parametric uncertainties
We consider models from mathematical neuroscience that are described by an ordinary differential equation (ODE) system of the form $\dot{\mathbf{x}}=f(\mathbf{x},Z)$, where $\mathbf{x}\in\mathbb{R}^{n}$ is the state vector and the uncertain system parameters of interest are stacked in a parametric uncertainty vector $Z\in\mathbb{R}^{d}$. We assume that the components of $Z$ are independent random variables, characterised by a known probability distribution supported on a bounded set representing the parameter space of interest for the considered model. Many different distributions of $Z$ may be handled by our proposed approach [74]. Since the uniform distribution has been observed to be the worst-case distribution for PRA [5], and we can theoretically transform any continuous random variable into a uniformly distributed one through a quantile transformation [32], in the absence of additional information on $Z$, we assume that each component $Z_{i}$, $i=1,\ldots,d$, is uniformly distributed over an uncertainty interval containing the nominal value $Z_{i}^{*}$. Precisely, we therefore have $Z_{i}\sim\mathcal{U}([Z_{i,{\min}},Z_{i,{\max}}])$, with $Z_{i,{\min}}\leq Z_{i}^{*}\leq Z_{i,{\max}}$, $i=1,\ldots,d$. The resulting ODE model subject to parametric uncertainty is
@@
where $f\colon\mathbb{R}^{n}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$, the hyper-rectangle $\mathcal{A}$ is the uncertainty bounding set and $\mathbf{x}(t;Z)$ is a stochastic process (Figure 1a); in this probabilistic setting, different system evolutions are possible depending on the specific realisation of the parameters in $Z$. In (1), stochasticity arises from probabilistic uncertainty in the system parameters, and not from a stochastic process driving the dynamics as it happens for instance in stochastic differential equations [51].
@@
To demonstrate our proposed methodology, we focus on two celebrated models from neuroscience: the Hindmarsh-Rose (HR) model for single-neuron dynamics [26] and the neural mass mean-field Jansen-Rit model for a single cortical column [30].
The HR model [26] is widely used to reproduce the spiking and bursting dynamics of single neurons observed in experimental data [23]; however, its robustness to probabilistic parametric uncertainty has not been studied so far. In the adimensionalised ODEs of the HR model
@@
variable $x_{1}(t)$ represents the neuron membrane potential, $x_{2}(t)$ a fast recovery current, and $x_{3}(t)$ a slow adaptation current that enables bursting (i.e., trains of multiple membrane potential spikes). $I$ is an injected current, $x_{R}$ is the reference resting membrane potential, $r$ is the timescale ratio between fast and slow currents, while parameters $a$, $b$, $c$, $d$, $s$ are related to the system nullclines [26]. We consider $Z_{\text{HR}}=[b,I]^{\top}$ as the parameters of interest, and set the other parameters to the standard values in [26]: $a=1$, $c=1$, $d=5$, $s=4$, $x_{R}=-8/5$, $r=0.01$. In a deterministic setting, for different values of the system parameters, the state trajectories $\mathbf{x}(t)=[x_{1}(t),x_{2}(t),x_{3}(t)]^{\top}$, $t\geq 0$, of this relatively simple model can exhibit multiple qualitatively distinct behaviours, which we call dynamical regimes. We focus on the membrane potential, $y(t)=x_{1}(t)$, as an output signal of interest; examples of realisations of $x_{1}(t)$ for different parameter values that pertain to different dynamical regimes, ranging from quiescence to bursting, are shown in Figure 2 (top).
The JR model [30] is a neural mass mean-field model that describes, rather than the dynamics of individual neurons, the generation of EEG signals by a single cortical column. In particular, the cortical column is represented through the interaction of the mean-field activity of three neuronal populations: pyramidal neurons, excitatory interneurons and inhibitory interneurons. In the corresponding ODE system
@@
the states $x_{i}(t)$, $i=1,2,3$, are neural potentials ($x_{1}$ is the mean excitatory postsynaptic potential, $x_{2}$ is the mean membrane potential of pyramidal cells, $x_{3}$ is the mean inhibitory postsynaptic potential), while $S(V)=\nu_{M}/\left(1+e^{r(V_{0}-V)}\right)$ is a sigmoidal function, where $\nu_{M}$ is the maximal firing rate of the family of neurons, $V_{0}$ is the excitability threshold of the population, $r$ is the firing threshold variance (i.e., the slope of the sigmoid at $V_{0}$). Parameters $A$ and $B$ are the maximum amplitude of excitatory and inhibitory postsynaptic potentials, respectively, while $a$ and $b$ are lumped parameters related to time constants in the dendritic network for excitatory and inhibitory synapses, respectively. All connectivity constants $C_{i}$, describing the strength of the links between pyramidal neurons and inhibitory and excitatory interneurons, are proportional to a single parameter $C$: $C_{1}=C$, $C_{2}=0.8C$, $C_{3}=0.25C$, $C_{4}=0.25C$, with $C=135$ [30]. We investigate $Z_{\text{JR}}=[A,B]$ as the parameters of interest, for varying values of $p$ and $C$, while all other parameters are fixed as in [30]: $a=100$, $b=50$, $V_{0}=6$, $\nu_{M}=5$, $r=0.56$. In a deterministic setting, the EEG-like output signal $y(t)=x_{2}(t)-x_{3}(t)$ exhibits different qualitative dynamical behaviours (regimes) as shown in Supplementary Figure S2.
@@
We consider the HR parameter space $b_{\min},b_{\max}]\times[I_{\min},I_{\max}]=[2.5,3.3]\times[2.2,4.4]$. Figure 1f shows the corresponding bifurcation diagram, from [7], in the deterministic setting: regions with different colours correspond to deterministic parameter values that result in regimes with different characteristics (exemplified in Figure 2). In our numerical experiments, we choose nominal parameter values $b^{*},I^{*}]^{\top}$ that lie within the interior of a region associated with one of the deterministic nominal regimes in Figure 1f. Probabilistic parametric uncertainty affecting the nominal parameter values may lead to loss of regime identifiability from the mean output trajectory realisation (see Figure 1c and Supplementary Figure S4 for examples). Our goal is thus to develop a computational framework that allows us to detect regimes in the presence of parametric uncertainty and identify the highest possible uncertainty level (corresponding to the largest possible neighbourhood of the nominal parameters, given predefined specifications) where the nominal regime, exhibited by the system output under the deterministic nominal parameters, is preserved in expectation. To this aim, we design dedicated metrics and apply them to the mean output signal, so as to relate it to the nominal regime.